# Enhancing Whole Slide Pathology Foundation Models through Stain Normalization

### v1

@@ Abstract
Recent advancements in digital pathology have led to the development of numerous foundational models that utilize self-supervised learning on patches extracted from gigapixel whole slide images (WSIs). While this approach leverages vast amounts of unlabeled data, we have discovered a significant issue: features extracted from these self-supervised models tend to cluster by individual WSIs, a phenomenon we term WSI-specific feature collapse. This problem can potentially limit the model’s generalization ability and performance on various downstream tasks. To address this issue, we introduce EXAPath , a novel foundational model trained on patches that have undergone stain normalization. Stain normalization helps reduce color variability arising from different laboratories and scanners, enabling the model to learn more consistent features. EXAPath is trained using 285,153,903 patches extracted from a total of 34,795 WSIs, combining data from The Cancer Genome Atlas (TCGA) and the Genotype-Tissue Expression (GTEx) project. Our experiments demonstrate that EXAPath significantly mitigates the feature collapse problem, indicating that the model has learned more generalized features rather than overfitting to individual WSI characteristics. We compared EXAPath with state-of-the-art models across six downstream task datasets, and our results show that EXAPath achieves excellent performance relative to the number of WSIs used and the model’s parameter count. This suggests that the application of stain normalization has substantially improved the model’s efficiency and generalization capabilities.
@@ 1 Introduction
In modern pathology, the analysis of digital pathology images plays a crucial role in cancer subtyping (Lu et al., 2021; Shao et al., 2021), prognosis prediction (Mobadersany et al., 2018; Chen et al., 2022b; Shmatko et al., 2022), and quantifying the tissue microenvironment (Schapiro et al., 2017; Mahmood et al., 2018; Saltz et al., 2018a; Graham et al., 2019). Whole Slide Images (WSIs), in particular, are high-resolution digitized tissue slides that often measure tens of thousands by tens of thousands of pixels in size. Due to the vast dimensions of these WSIs, many researchers have adopted the approach of dividing the images into smaller patches to train their models. A single WSI can be divided into thousands or even tens of thousands of patches, but accurately labeling each of these patches is a significant burden in terms of time and cost. To overcome this limitation, recent approaches have widely adopted self-supervised learning techniques to first train a foundation model, which is then fine-tuned for various downstream tasks (Kang et al., 2023; Chen et al., 2024; Vorontsov et al., 2023; Dippel et al., 2024). This approach allows for the effective utilization of large amounts of unlabeled data while achieving high performance with limited labeled data.
@@
In this paper, we first introduce our novel observation that the features of models trained through self-supervised learning exhibit an unexpectedly high degree of overlap. Specifically, we divide whole slide images into patches and create a foundation model using the DINO (self-DIstillation with NO labels) (Caron et al., 2021) self-supervised learning method. During this process, as the whole slides are divided into small patches for training, the model has no information about which whole slide each specific patch originated from. Subsequently, we extract features of each patch and visualize them using t-SNE (van der Maaten & Hinton, 2008). Surprisingly, we observe that the patch features cluster according to their respective WSIs. We refer to this phenomenon as WSI-specific feature collapse. Whole slides undergo Hematoxylin and Eosin (H&E) staining to clearly reveal tissue structures. We hypothesize that this clustering phenomenon occurs due to variations in staining intensity between laboratories or even within the same laboratory, resulting in color differences among WSIs. In whole slide image analysis, model features should focus on capturing pathologically significant information rather than simply relying on the degree of staining. Critical information that should be observed includes nuclear size and shape, cell density, and structural changes in tissue. If the model concentrates excessively on staining variations, it may overlook these essential pathological features, potentially leading to inaccuracies in important downstream tasks such as cancer type classification, mutation classification, and survival analysis. Therefore, it is crucial to ensure that the model is trained to properly recognize and learn these important pathological characteristics during the training process. To address this issue, we first apply Macenko normalization (Macenko et al., 2009) to the patches before feature extraction in an effort to standardize the color characteristics across patches. While this somewhat mitigates the feature collapse, significant collapse is still observed. To further address this challenge, we introduce EXAPath, a stain-normalized pathology foundation model. This model is created by performing DINO self-supervised learning pretraining on patches that have undergone Macenko normalization. EXAPath is trained using 285,153,903 patches extracted from a total of 34,795 WSIs, combining data from The Cancer Genome Atlas (TCGA) (Weinstein et al., 2013) and the Genotype-Tissue Expression (GTEx) project (Lonsdale et al., 2013). We have confirmed that EXAPath significantly mitigates the feature collapse problem. This implies that the model has a higher potential to extract pathologically relevant features crucial for downstream tasks, rather than simply relying on color-based characteristics. These improvements are expected to enhance the model’s performance and reliability across various pathological analysis tasks. We validated the performance of EXAPath on various patch-level tasks including PCAM (Veeling et al., 2018; Bejnordi et al., 2017), MHIST (Wei et al., 2021), CRC 100k (Kather et al., 2018), TIL-Detection (Abousamra et al., 2022; Saltz et al., 2018b; Kaczmarzyk et al., ), MSI-CRC, and MSI-STAD (Kather, 2019) dataset. Notably, despite being trained on relatively less data compared to existing State-of-the-Art (SOTA) models, and using only publicly available datasets, EXAPath achieved comparable performance levels to these SOTA models across these diverse tasks. Our contributions can be summarized as follows:
@@
Figure 1: Performance comparison of models based on the number of parameters and the number of WSIs used for training. The average Top-1 accuracy represents the mean linear evaluation performance across six downstream tasks. (a) Average Top-1 accuracy versus the number of parameters. (b) Average Top-1 accuracy versus the number of WSIs used for training. Notably, our model (EXAPath) achieves high performance despite having fewer parameters and using fewer WSIs compared to other models, demonstrating its efficiency.
@@
  • To address this issue, we develop EXAPath, a novel model that applies stain normalization to patches during foundation model training. We demonstrate that EXAPath effectively mitigates the feature collapse problem.
  • We evaluate EXAPath against state-of-the-art models on six diverse patch-level tasks. The results show that EXAPath exhibits superior performance relative to the number of WSIs used and the number of model parameters.
@@ 2 Related Work
Pathology Image Foundation Models. With the significant advancements in self-supervised learning (SSL) methods, there is a notable trend in applying these methods, commonly used in natural image processing, to histopathology to develop robust foundation models for whole slide images (WSIs) (Gildenblat & Klaiman, 2019; Ciga et al., 2022; Yang et al., 2022; Kang et al., 2023). Given the gigapixel resolution of WSIs, creating foundation models by dividing images into patch units has become a prevalent approach. For example, REMEDIS (Azizi et al., 2023) employs SimCLR (Chen et al., 2020a) to train on various medical domain images, including pathology. Similarly, Ciga et al. (2022) utilizes SimCLR for training on histopathology images. Kang et al. (2023) applies four SSL methods—MoCo v2 (Chen et al., 2020b), SwAV (Caron et al., 2020), Barlow Twins (Zbontar et al., 2021), and DINO (Caron et al., 2021)—to pathology data and evaluates their performance on image classification and nuclei instance segmentation tasks. Campanella et al. (2023) compares the Masked Autoencoder (MAE) (He et al., 2022) and DINO (Caron et al., 2021) algorithms using the largest pathology dataset to date, which comprises over three billion images. Their evaluation across six clinical tasks demonstrates that pre-training on pathology data is more effective than pre-training on natural images, with the DINO algorithm showing superior generalization performance across all tasks. HIPT (Chen et al., 2022a) also implements DINO within a hierarchical structure. Phikon (Filiot et al., 2023) employs iBOT (Zhou et al., 2022), a self-supervised learning method based on masked image modeling (MIM), for histology images. The ViT-Base model using iBOT exhibits excellent performance across 17 downstream tasks on 7 cancer types, suggesting its potential for developing a foundation model for histopathology. CTransPath (Wang et al., 2022) introduces a new self-supervised learning strategy called Semantically-Relevant Contrastive Learning (SRCL), an enhancement of MoCo v3 (Chen* et al., 2021). Additionally, UNI (Chen et al., 2024), RudolfV (Dippel et al., 2024), Virchow (Vorontsov et al., 2023), and Prov-GigaPath (Xu et al., 2024) utilize DINO v2 for training pathology foundation models. Prov-GigaPath (Xu et al., 2024) applies the LongNet (Ding et al., 2023) architecture to learn slide-level representations.
@@
Early methods focused on histogram transformation, such as Reinhard normalization (Reinhard et al., 2001), which employs statistical techniques to transfer the color characteristics from one image to another by converting them to the Lab color space. This method is widely recognized for its simplicity and effectiveness in standardizing color distributions across different images. Another notable approach is Macenko normalization (Macenko et al., 2009), which transforms images into an optical density (OD) space. PCA is then performed on the OD space to identify the principal components corresponding to the Hematoxylin and Eosin stains, thus enabling effective stain separation and normalization. Inspired by the biological fact about stain binding, Vahadane et al. (2016) extended the use of nonnegative matrix factorization (NMF) (Rabinovich et al., 2003) by introducing sparsity constraints, which enhance the robustness of stain separation by limiting the number of non-zero components, thereby aligning with the actual biological composition of histological samples.
Recently, deep learning-based approaches have advanced stain normalization techniques, particularly through the use of image-to-image translation architectures combined with adversarial learning methods (Bentaieb & Hamarneh, 2018). These approaches leverage generative adversarial networks (GANs) (Goodfellow et al., 2014), such as CycleGAN (Zhu et al., 2017) and StarGAN (Choi et al., 2018), to learn the mappings between different staining styles. By employing adversarial training, these models are able to produce normalized images that closely resemble the target staining characteristics while preserving essential structural features of the tissue (Wagner et al., 2021; Nazki et al., 2023). Such advancements demonstrate the potential of deep learning to overcome the limitations of traditional stain normalization methods, offering more robust and scalable solutions for digital pathology.
@@ 3 Self-supervised Pre-training for Pathology
This section introduces the ‘WSI-specific feature collapse’ that occurs when pathology foundation models are trained using self-supervised learning, and discusses methods to reduce this phenomenon. In Section 3.1, we discuss the ’WSI-specific feature collapse’ problem. This refers to the phenomenon where features from the same WSI tend to cluster together, even though each patch is learned without information about which WSI it comes from during the training process. Section 3.2 demonstrates that applying stain normalization before feature extraction does not significantly reduce the collapse. In Section 3.3, we introduce EXAPath, which applies stain normalization to the training data from the beginning of foundation model training to further reduce ‘WSI-specific feature collapse’.
@@
DINO (Caron et al., 2021) involves applying different augmentations to an image and inputting them into both teacher and student models, training them to produce identical features. Color jittering is a data augmentation technique commonly used in self-supervised learning, including DINO. It involves randomly altering the colors of an image to create different variations, helping the model learn robust features invariant to color changes. However, despite the use of color jittering augmentation, when visualizing patch features using t-SNE (van der Maaten & Hinton, 2008), we observe that features still tend to collapse according to their source WSI. Specifically, we trained a foundation model using DINO on $256^{2}$-sized patches extracted from gigapixel WSIs. It is important to note that the neural network is trained only on these isolated patches, without any information about which WSI each patch originated from. We randomly selected 1,000 patches from each of 10 arbitrary WSIs, performed a forward pass through the trained foundation model, and then visualized the resulting 10,000 features using t-SNE (see Figure 2(a)). Interestingly, even though the model does not know the source WSI of each patch, we observe that patches from the same WSI tend to have similar features and cluster together. This phenomenon suggests that, instead of learning features crucial for downstream tasks and pathologically significant, the model may have unintentionally learned WSI-specific characteristics, such as staining intensity or subtle differences due to scanner variations. This feature collapse phenomenon could potentially compromise the model’s ability to generalize. If the model becomes overly adapted to individual WSI characteristics, its performance may degrade when applied to new WSIs or images acquired from different institutions. This result suggests that when creating a pretrained foundation model for WSIs, rather than directly applying self-supervised learning methods used for natural images, we need to use approaches that take into account the specific characteristics of WSIs.
@@
We hypothesized that the WSI-specific feature collapse occurs due to variations in staining intensity across different WSIs, which result in color differences. This phenomenon arises from the staining process of WSIs, which depends on the types of stains used and the degree of staining. Additionally, differences in slide scanner settings or performance can affect color variations in digitized WSIs. Therefore, to address these color differences, we employed Macenko stain normalization (Macenko et al., 2009) before extracting features. We used the same pretrained model and patches as described in Section 3.1. Figure 2(b) shows the results. Compared to Figure 2(a), where features are completely separated by WSI, Figure 2(b) shows that features from different WSIs are now slightly mixed, but they still tend to cluster by WSI. This indicates that while Macenko normalization partially mitigates color differences between WSIs, it is not sufficient to fully resolve the issue, as features from the same WSI still exhibit noticeable clustering and the collapse problem remains significant. Furthermore, it suggests that applying Macenko normalization only during inference has limited effectiveness because the pretrained model has already learned to extract features biased toward color characteristics.
@@
We use DINO (Caron et al., 2021) self-supervised learning to train a WSI foundation model. We refer to this model as EXAPath, which is trained with stain-normalized data using DINO. DINO involves applying different augmentations to an image and inputting them into both teacher and student models, training them to produce identical features. The teacher model, updated with an exponential moving average (EMA) of the student parameters, guides the output of the student model. The student model is trained using cross-entropy loss to measure their similarity. By inputting differently augmented images into the teacher and student models, the features produced by both models are compared and aligned. In the original DINO paper, different augmentation methods are applied to two $256^{2}$ global views and several $96^{2}$ local views of each image.
Combining Macenko normalization with the existing DINO augmentation is a design choice. We apply Macenko normalization to all images with 100% probability before applying DINO augmentation. This approach is chosen because applying Macenko normalization to each global view and local view individually is computationally intensive, creating a bottleneck in the data loader during training. Figure 2(c) visualizes the features of EXAPath trained with DINO using Macenko-normalized images. Macenko normalization is applied not only during training but also when extracting features for each patch in the visualization process. Compared to Figures 2(a) and 2(b), we observe that patch features from various WSIs are much more intermingled. This indicates that the model learns more generalized features and is less dependent on the source of the WSIs. This allows the model to focus on more histologically relevant features. We still observe that some clusters of features are grouped by WSI. We believe this occurs for two reasons: first, while Macenko normalization significantly reduces color differences, it may not completely eliminate them; second, patches from the same WSI likely share clinically relevant features beyond just color, due to variations in tissue type, cancer subtype, and patient characteristics within each WSI.

### v3

# EXAONEPath 1.0 Patch-level Foundation Model for Pathology

@@ Abstract
Recent advancements in digital pathology have led to the development of numerous foundational models that utilize self-supervised learning on patches extracted from gigapixel whole slide images (WSIs). While this approach leverages vast amounts of unlabeled data, we have discovered a significant issue: features extracted from these self-supervised models tend to cluster by individual WSIs, a phenomenon we term WSI-specific feature collapse. This problem can potentially limit the model’s generalization ability and performance on various downstream tasks. To address this issue, we introduce EXAONEPath111github:https://github.com/LG-AI-EXAONE/EXAONEPath
huggingface:https://huggingface.co/LGAI-EXAONE/EXAONEPath
, a novel foundational model trained on patches that have undergone stain normalization. Stain normalization helps reduce color variability arising from different laboratories and scanners, enabling the model to learn more consistent features. EXAONEPath is trained using 285,153,903 patches extracted from a total of 34,795 WSIs. Our experiments demonstrate that EXAONEPath significantly mitigates the feature collapse problem, indicating that the model has learned more generalized features rather than overfitting to individual WSI characteristics. We compared EXAONEPath with state-of-the-art models across six downstream task datasets, and our results show that EXAONEPath achieves superior performance relative to the number of WSIs used and the model’s parameter count. This suggests that the application of stain normalization has substantially improved the model’s efficiency and generalization capabilities.
*Important Notice*
This document is valid only for EXAONEPath version 1.0. Also, please refer to the license at the end of the document, when using this model.
@@ 1 Introduction
In modern pathology, the analysis of digital pathology images plays a crucial role in cancer subtyping [Lu et al., 2021, Shao et al., 2021], prognosis prediction [Mobadersany et al., 2018, Chen et al., 2022b, Shmatko et al., 2022], and quantifying the tissue microenvironment [Schapiro et al., 2017, Mahmood et al., 2018, Saltz et al., 2018a, Graham et al., 2019]. Whole Slide Images (WSIs), in particular, are high-resolution digitized tissue slides that often measure tens of thousands by tens of thousands of pixels in size. Due to the vast dimensions of these WSIs, many researchers have adopted the approach of dividing the images into smaller patches to train their models. A single WSI can be divided into thousands or even tens of thousands of patches, but accurately labeling each of these patches is a significant burden in terms of time and cost. To overcome this limitation, recent approaches have widely adopted self-supervised learning techniques to first train a foundation model, which is then fine-tuned for various downstream tasks [Kang et al., 2023, Chen et al., 2024, Vorontsov et al., 2023, Dippel et al., 2024]. This approach allows for the effective utilization of large amounts of unlabeled data while achieving high performance with limited labeled data.
@@
In this paper, we first introduce our novel observation that the features of models trained through self-supervised learning exhibit an unexpectedly high degree of overlap. Specifically, we divide whole slide images into patches and create a foundation model using the DINO (self-DIstillation with NO labels) [Caron et al., 2021] self-supervised learning method. During this process, as the whole slides are divided into small patches for training, the model has no information about which whole slide each specific patch originated from. Subsequently, we extract features of each patch and visualize them using t-SNE [van der Maaten and Hinton, 2008]. Surprisingly, we observe that the patch features cluster according to their respective WSIs. We refer to this phenomenon as WSI-specific feature collapse. Whole slides undergo Hematoxylin and Eosin (H&E) staining to clearly reveal tissue structures. We hypothesize that this clustering phenomenon occurs due to variations in staining intensity between laboratories or even within the same laboratory, resulting in color differences among WSIs. In whole slide image analysis, model features should focus on capturing pathologically significant information rather than simply relying on the degree of staining. Critical information that should be observed includes nuclear size and shape, cell density, and structural changes in tissue. If the model concentrates excessively on staining variations, it may overlook these essential pathological features, potentially leading to inaccuracies in important downstream tasks such as cancer type classification, mutation classification, and survival analysis. Therefore, it is crucial to ensure that the model is trained to properly recognize and learn these important pathological characteristics during the training process. To address this issue, we first apply Macenko normalization [Macenko et al., 2009] to the patches before feature extraction in an effort to standardize the color characteristics across patches. While this somewhat mitigates the feature collapse, significant collapse is still observed. To further address this challenge, we introduce EXAONEPath, a stain-normalized pathology foundation model. This model is created by performing DINO self-supervised learning pretraining on patches that have undergone Macenko normalization. EXAONEPath is trained using 285,153,903 patches extracted from a total of 34,795 WSIs. We have confirmed that EXAONEPath significantly mitigates the feature collapse problem. This implies that the model has a higher potential to extract pathologically relevant features crucial for downstream tasks, rather than simply relying on color-based characteristics. These improvements are expected to enhance the model’s performance and reliability across various pathological analysis tasks. We validated the performance of EXAONEPath on various patch-level tasks including PCAM [Veeling et al., 2018, Bejnordi et al., 2017], MHIST [Wei et al., 2021], CRC 100k [Kather et al., 2018], TIL-Detection [Abousamra et al., 2022, Saltz et al., 2018b, Kaczmarzyk et al., ], MSI-CRC, and MSI-STAD [Kather, 2019] dataset. Notably, despite being trained on relatively less data compared to existing State-of-the-Art (SOTA) models, and using only publicly available datasets, EXAONEPath achieved comparable performance levels to these SOTA models across these diverse tasks. Our contributions can be summarized as follows:
@@
Figure 1: Performance comparison of models based on the number of parameters and the number of WSIs used for training. The average Top-1 accuracy represents the mean linear evaluation performance across six downstream tasks. (a) Average Top-1 accuracy versus the number of parameters. (b) Average Top-1 accuracy versus the number of WSIs used for training. Notably, our model (EXAONEPath) achieves high performance despite having fewer parameters and using fewer WSIs compared to other models, demonstrating its efficiency.
@@
  • To address this issue, we develop EXAONEPath, a novel model that applies stain normalization to patches during foundation model training. We demonstrate that EXAONEPath effectively mitigates the feature collapse problem.
  • We evaluate EXAONEPath against state-of-the-art models on six diverse patch-level tasks. The results show that EXAONEPath exhibits superior performance relative to the number of WSIs used and the number of model parameters.
@@ 2 Related Work
Pathology Image Foundation Models. With the significant advancements in self-supervised learning (SSL) methods, there is a notable trend in applying these methods, commonly used in natural image processing, to histopathology to develop robust foundation models for whole slide images (WSIs) [Gildenblat and Klaiman, 2019, Ciga et al., 2022, Yang et al., 2022, Kang et al., 2023]. Given the gigapixel resolution of WSIs, creating foundation models by dividing images into patch units has become a prevalent approach. For example, REMEDIS [Azizi et al., 2023] employs SimCLR [Chen et al., 2020a] to train on various medical domain images, including pathology. Similarly, Ciga et al. [2022] utilizes SimCLR for training on histopathology images. Kang et al. [2023] applies four SSL methods—MoCo v2 [Chen et al., 2020b], SwAV [Caron et al., 2020], Barlow Twins [Zbontar et al., 2021], and DINO [Caron et al., 2021]—to pathology data and evaluates their performance on image classification and nuclei instance segmentation tasks. Campanella et al. [2023] compares the Masked Autoencoder (MAE) [He et al., 2022] and DINO [Caron et al., 2021] algorithms using the largest pathology dataset to date, which comprises over three billion images. Their evaluation across six clinical tasks demonstrates that pre-training on pathology data is more effective than pre-training on natural images, with the DINO algorithm showing superior generalization performance across all tasks. HIPT [Chen et al., 2022a] also implements DINO within a hierarchical structure. Phikon [Filiot et al., 2023] employs iBOT [Zhou et al., 2022], a self-supervised learning method based on masked image modeling (MIM), for histology images. The ViT-Base model using iBOT exhibits excellent performance across 17 downstream tasks on 7 cancer types, suggesting its potential for developing a foundation model for histopathology. CTransPath [Wang et al., 2022] introduces a new self-supervised learning strategy called Semantically-Relevant Contrastive Learning (SRCL), an enhancement of MoCo v3 [Chen* et al., 2021]. Additionally, UNI [Chen et al., 2024], RudolfV [Dippel et al., 2024], Virchow [Vorontsov et al., 2023], and Prov-GigaPath [Xu et al., 2024] utilize DINO v2 for training pathology foundation models. Prov-GigaPath [Xu et al., 2024] applies the LongNet [Ding et al., 2023] architecture to learn slide-level representations.
@@
Early methods focused on histogram transformation, such as Reinhard normalization [Reinhard et al., 2001], which employs statistical techniques to transfer the color characteristics from one image to another by converting them to the Lab color space. This method is widely recognized for its simplicity and effectiveness in standardizing color distributions across different images. Another notable approach is Macenko normalization [Macenko et al., 2009], which transforms images into an optical density (OD) space. PCA is then performed on the OD space to identify the principal components corresponding to the Hematoxylin and Eosin stains, thus enabling effective stain separation and normalization. Inspired by the biological fact about stain binding, Vahadane et al. [2016] extended the use of nonnegative matrix factorization (NMF) [Rabinovich et al., 2003] by introducing sparsity constraints, which enhance the robustness of stain separation by limiting the number of non-zero components, thereby aligning with the actual biological composition of histological samples.
Recently, deep learning-based approaches have advanced stain normalization techniques, particularly through the use of image-to-image translation architectures combined with adversarial learning methods [Bentaieb and Hamarneh, 2018]. These approaches leverage generative adversarial networks (GANs) [Goodfellow et al., 2014], such as CycleGAN [Zhu et al., 2017] and StarGAN [Choi et al., 2018], to learn the mappings between different staining styles. By employing adversarial training, these models are able to produce normalized images that closely resemble the target staining characteristics while preserving essential structural features of the tissue [Wagner et al., 2021, Nazki et al., 2023]. Such advancements demonstrate the potential of deep learning to overcome the limitations of traditional stain normalization methods, offering more robust and scalable solutions for digital pathology.
@@ 3 Self-supervised Pre-training for Pathology
This section introduces the ‘WSI-specific feature collapse’ that occurs when pathology foundation models are trained using self-supervised learning, and discusses methods to reduce this phenomenon. In Section 3.1, we discuss the ’WSI-specific feature collapse’ problem. This refers to the phenomenon where features from the same WSI tend to cluster together, even though each patch is learned without information about which WSI it comes from during the training process. Section 3.2 demonstrates that applying stain normalization before feature extraction does not significantly reduce the collapse. In Section 3.3, we introduce EXAONEPath, which applies stain normalization to the training data from the beginning of foundation model training to further reduce ‘WSI-specific feature collapse’.
@@
DINO [Caron et al., 2021] involves applying different augmentations to an image and inputting them into both teacher and student models, training them to produce identical features. Color jittering is a data augmentation technique commonly used in self-supervised learning, including DINO. It involves randomly altering the colors of an image to create different variations, helping the model learn robust features invariant to color changes. However, despite the use of color jittering augmentation, when visualizing patch features using t-SNE [van der Maaten and Hinton, 2008], we observe that features still tend to collapse according to their source WSI. Specifically, we trained a foundation model using DINO on $256^{2}$-sized patches extracted from gigapixel WSIs. It is important to note that the neural network is trained only on these isolated patches, without any information about which WSI each patch originated from. We randomly selected 1,000 patches from each of 10 arbitrary WSIs, performed a forward pass through the trained foundation model, and then visualized the resulting 10,000 features using t-SNE (see Figure 2(a)). Interestingly, even though the model does not know the source WSI of each patch, we observe that patches from the same WSI tend to have similar features and cluster together. This phenomenon suggests that, instead of learning features crucial for downstream tasks and pathologically significant, the model may have unintentionally learned WSI-specific characteristics, such as staining intensity or subtle differences due to scanner variations. This feature collapse phenomenon could potentially compromise the model’s ability to generalize. If the model becomes overly adapted to individual WSI characteristics, its performance may degrade when applied to new WSIs or images acquired from different institutions. This result suggests that when creating a pretrained foundation model for WSIs, rather than directly applying self-supervised learning methods used for natural images, we need to use approaches that take into account the specific characteristics of WSIs.
@@
We hypothesized that the WSI-specific feature collapse occurs due to variations in staining intensity across different WSIs, which result in color differences. This phenomenon arises from the staining process of WSIs, which depends on the types of stains used and the degree of staining. Additionally, differences in slide scanner settings or performance can affect color variations in digitized WSIs. Therefore, to address these color differences, we employed Macenko stain normalization [Macenko et al., 2009] before extracting features. We used the same pretrained model and patches as described in Section 3.1. Figure 2(b) shows the results. Compared to Figure 2(a), where features are completely separated by WSI, Figure 2(b) shows that features from different WSIs are now slightly mixed, but they still tend to cluster by WSI. This indicates that while Macenko normalization partially mitigates color differences between WSIs, it is not sufficient to fully resolve the issue, as features from the same WSI still exhibit noticeable clustering and the collapse problem remains significant. Furthermore, it suggests that applying Macenko normalization only during inference has limited effectiveness because the pretrained model has already learned to extract features biased toward color characteristics.
@@
We use DINO [Caron et al., 2021] self-supervised learning to train a WSI foundation model. We refer to this model as EXAONEPath, which is trained with stain-normalized data using DINO. DINO involves applying different augmentations to an image and inputting them into both teacher and student models, training them to produce identical features. The teacher model, updated with an exponential moving average (EMA) of the student parameters, guides the output of the student model. The student model is trained using cross-entropy loss to measure their similarity. By inputting differently augmented images into the teacher and student models, the features produced by both models are compared and aligned. In the original DINO paper, different augmentation methods are applied to two $256^{2}$ global views and several $96^{2}$ local views of each image.
Combining Macenko normalization with the existing DINO augmentation is a design choice. We apply Macenko normalization to all images with 100% probability before applying DINO augmentation. This approach is chosen because applying Macenko normalization to each global view and local view individually is computationally intensive, creating a bottleneck in the data loader during training. Figure 2(c) visualizes the features of EXAONEPath trained with DINO using Macenko-normalized images. Macenko normalization is applied not only during training but also when extracting features for each patch in the visualization process. Compared to Figures 2(a) and 2(b), we observe that patch features from various WSIs are much more intermingled. This indicates that the model learns more generalized features and is less dependent on the source of the WSIs. This allows the model to focus on more histologically relevant features. We still observe that some clusters of features are grouped by WSI. We believe this occurs for two reasons: first, while Macenko normalization significantly reduces color differences, it may not completely eliminate them; second, patches from the same WSI likely share clinically relevant features beyond just color, due to variations in tissue type, cancer subtype, and patient characteristics within each WSI.